# Bias Amplification: Language Models as Increasingly Biased Media

### v1

@@ 8 Ethical Considerations
This study addresses bias amplification in LLMs, a technical phenomenon with profound ethical implications, particularly regarding fairness and the integrity of AI systems. The risk of bias amplification is especially concerning in systems that are iteratively trained on synthetic data, as it can lead to unintended distortions in model outputs. These distortions may propagate harmful biases, influencing downstream tasks in areas such as automated content generation, decision-making, and user interactions with AI.
@@
Moreover, the identification of distinct neural mechanisms for bias amplification and model collapse brings to light the challenges of ensuring equitable performance across all dimensions of model behavior and raises important ethical questions regarding the adequacy of current mitigation techniques, particularly in high-stakes scenarios where the cost of algorithmic bias is substantial.
Future research should prioritize the development of more comprehensive and domain-specific bias mitigation techniques, with a clear focus on minimizing ethical risks. Additionally, rigorous testing and validation across diverse datasets and real-world applications will be critical to ensuring that models trained using these methods do not exacerbate existing inequalities or produce harmful outcomes.
@@ Appendix A Mathematical Formulation of WMLE
Appendix B Fine-tuning Setup
@@
Appendix C Qualitative Bias Analysis Framework and Example of Bias Amplification Across Generations
We employed qualitative methods to confirm our findings in media bias. Specifically, we utilized a media bias identification framework grounded in foundational works such as Entman’s framing theory (Entman, 1993) and other research on media bias detection (Rodrigo-Ginés et al., 2024; Groeling, 2013). This framework provides a robust lens to evaluate political biases in the framing and language use of media texts (explained below). Given the nature of our data—text exclusive of visual or contextual cues like formatting—certain types of media bias commonly seen in formatted articles or televised programs (e.g., visual bias or tone) may not apply. Therefore, our focus was on the two key aspects of political bias that are particularly relevant in textual analysis:
@@ Appendix C Qualitative Bias Analysis Framework and Example of Bias Amplification Across Generations
    Analysis of Loaded Language Bia:
@@
Appendix D Distribution of Text Quality Index Across Generations
D.1 Deterministic Setup
Figure 8: This figure presents the distribution of the text quality index across multiple generations of synthetic articles, for the deterministic setup of GPT-2.
D.2 Beam Search Setup
Figure 9: This figure presents the distribution of text quality index across multiple generations of synthetic articles, for the beam search setup of GPT-2.
D.3 Nucleus Sampling Setup
Figure 10: This figure presents the distribution of text quality index across multiple generations of synthetic articles, for the nucleus sampling setup of GPT-2.
Appendix E Average Perplexity Across Generations
Figure 11: This figure shows the average perplexity for synthetic articles , along with 95% confidence intervals, across multiple generations for three generation setups, with ’synthetic’ representing the determistic setup
Appendix F Example of Quality Deterioration Across Generations
@@
F.1 Example of Deterministic Setup
@@
  • Generation 1: , says US official, says US military is ready to go, says Pentagon official. Says US military is ready to go, says Pentagon official. Says US military is ready to go, says Pentagon official. Says US military is ready to go, says Pentagon official. Says US military is ready to go, says Pentagon official the world is not so simple. Text Quality Index: 2.44; Perplexity: 5.77
  • Generation 5: in Iraq and Syria (ISIS) group Read more The Iraqi army has been fighting the Islamic State since the group seized large swaths of territory in Iraq and Syria (ISIS) group in 2014. The Iraqi army has been fighting the Islamic State the Iraqi army. The move comes as the U.S. the Iraqi the the the the holiest places in the world. Text Quality Index: 2.01; Perplexity: 3.17
@@
F.2 Example of Beam Search Setup
  • Original Article: Senate approval will be a much higher bar to clear: There’s no shortage of bipartisan criticism from the upper chamber right now, and the GOP holds a considerably smaller majority margin on that half of the Hill. Text Quality Index: 2.70
  • Generation 0: on Thursday, with Senate Majority Leader Mitch McConnell (R-Ky.) announcing that the bill would be headed to the House for a vote. Senate Majority Leader Harry Reid (D-Nev.) said in a statement that the Senate will vote on the bill ""as soon as possible." Text Quality Index: 2.26; Perplexity: 7.99
  • Generation 1: Read more “It’s going to take a lot of work,” House Speaker John Boehner (R-Ohio) said on ABC’s “This Week.” “It’s going to take a lot of work,” House Speaker John Boehner (R-Ohio. “It’s going to take a lot of work,” said Sen. John McCain, R-Ariz., chairman of the Senate Armed Services Committee. “It’s going to take a lot of patience,” he added. “It’s going to take Recovery and Reinvestment Act of 2009.Text Quality Index: 2.26; Perplexity: 5.42
  • Generation 5: committee MORE (R-Ohio) and Senate Majority Leader Mitch McConnell Addison (Mitch) Mitchell McConnellSchumer: Nunes gave Trump’secretly altered’ version of memo Davis: ‘Deep state’ existed in “ House Speaker John A. Boehner House Speaker House Speaker House Speaker John A. Boehner of the House Speaker House Speaker House Speaker John A. Boehner (R-Ohio) and Senate Majority Leader Mitch McConnell Addison (Mitch) Mitchell McConnellSchumer: Nunes gave Trump’secretly altered’ version of memo Davis: ‘Deep state’ existed in ’16 – but it elected Trump Health Care’s “I don’t know. Text Quality Index: 1.07; Perplexity: 3.33
  • Generation 10: vote House Speaker John A. Boehner (R-Ohio) and Senate Majority Leader Mitch McConnell Addison (Mitch) Mitchell McConnellSchumer: Nunes gave Trump’secretly “ House Speaker John A. Boehner (R-Ohio) House Speaker House Speaker House of “ House Speaker House Speaker House Speaker House of Senate Senate Majority Senate Majority Senate Majority Senate Majority Senate Majority Senate Majority Senate Majority Leader Senate Majority Leader Mitch McConnell Addison (Mitch) Mitchell McConnellSchumer Text Quality Index: 0.93; Perplexity: 3.45
F.3 Example of Nucleus Sampling Setup
  • Original Article: Senate approval will be a much higher bar to clear: There’s no shortage of bipartisan criticism from the upper chamber right now, and the GOP holds a considerably smaller majority margin on that half of the Hill. Text Quality Index: 2.70
  • Generation 0: , but no major votes on amendments Republicans said would be needed after a day of debate. Senate GOP senators announced Thursday they were giving the bill to the House for a vote on the Senate floor before heading to Washington for a vote on the bill, according to reports. The vote is expected to be postponed amid a GOP filibuster is about to vote. Text Quality Index: 2.47; Perplexity: 19.13
  • Generation 1: , House, after GOP leaders failed to reach a deal on funding legislation The House voted Thursday to proceed on a measure to overhaul the nation’s health care system, with the final vote coming in the upper chamber. But Senate Republicans, angered by the failure of to reach a deal on funding bill, are expected to is now set to. Text Quality Index: 2.31; Perplexity: 14.21
  • Generation 5: , White House says MORE (R-Neb.) said Monday that President Obama’s health care law is on track to pass the Senate with bipartisan support. The White House on Tuesday signaled the vote was a vote of 216-189. Text Quality Index: 2.20; Perplexity: 18.64
  • Generation 10: vote Thursday afternoon, sending bill on track to head to President Trump for final vote Thursday afternoon, sending bill on track to head to President Trump for final vote Thursday afternoon, sending bill on track to. The bill to move forward with the bill. ""This is a. ""The the law is the. Text Quality Index: 1.59; Perplexity: 27.51
Appendix G Pearson Correlation Between Neuron Weight and Bias Performance
Figure 12: Scatter plot showing the Pearson correlation between neuron weight and bias performance (i.e. percentage of right-leaning articles). Each point represents a neuron’s correlation at a specific layer.
Appendix H Pearson Correlation Between Neuron Weight and Generation Quality
Figure 13: Scatter plot showing the Pearson correlation between neuron weight and generation quality (i.e. average text quality index). Each point represents a neuron’s correlation at a specific layer.
Appendix I Pearson Correlation Between Neuron Activation and Bias Performance
Figure 14: Scatter plot showing the Pearson correlation between neuron activation value and bias performance (i.e. percentage of right-leaning articles). Each point represents a neuron’s correlation at a specific layer.
Appendix J Pearson Correlation Between Neuron Activation and Generation Quality
Figure 15: Scatter plot showing the Pearson correlation between neuron activation value and generation quality (i.e. average text quality index). Each point represents a neuron’s correlation at a specific layer.
Appendix K Mathematical Details for the Statistical Tests
@@ Appendix I Pearson Correlation Between Neuron Activation and Bias Performance
We will explain how the relationship between changes in neuron weight and changes in bias performance can be statistically tested. The same approach applies to the other three relationships discussed in Section 5.4.
K.1 Notations
  • $x_{i,j}$ denotes the weight of neuron $j$ in model $i$, where $i$ ranges from 1 to 66 and $j$ ranges from 1 to 9216.
  • $y_{i}$ denotes the percentage of right-leaning articles generated by model $i$, where $i$ ranges from 1 to 66.
  • $\Delta x_{i,j}$ denotes the change in the weight of neuron $j$ from model $i-1$ to model $i$ under a specific setup.
  • $\Delta y_{i}$ denotes the change in the percentage of right-leaning articles generated from model $i$ relative to the percentage generated the model $i-1$.
K.2 Regression
The regression we use is:
$\Delta y_{i}=\alpha_{j}+\beta_{j}\Delta x_{i,j}+\epsilon_{i,j}$ (3)
where $\alpha_{j}$ is the constant term for neuron j, $\beta_{j}$ measures the linear relationship between $\Delta x_{i,j}$ and $\Delta y_{i}$, and $\epsilon_{i,j}$ is the residual of the regression. Here, we apply a first-order difference to $x_{i}$ and $y_{i}$ to eliminate potential serial correlations in $x_{i}$ and $y_{i}$.
K.3 Statistical Test
@@
Appendix L Literature Review of Model Collapse

### v2

@@ 8 Ethical Considerations
This study focus on bias amplification in LLMs, a phenomenon with profound ethical implications, particularly regarding fairness and the integrity of AI systems. The risk of bias amplification is especially concerning in systems that are iteratively trained on synthetic data, as it can lead to unintended distortions in model outputs. These distortions may propagate harmful biases, influencing downstream tasks in areas such as automated content generation, decision-making, and user interactions with AI.
@@
It is crucial to explicitly state that the methodologies and data used in this research should not be applied to develop or train biased models for harmful applications. This study aims to enhance the understanding of bias amplification and model collapse in LLMs while promoting responsible and ethical AI development.
This work includes content that may contain personally identifying information or offensive language. However, all such material is derived solely from publicly available news article datasets or generated synthetically by models fine-tuned on these open-source datasets—or on synthetic data produced by earlier generations in our training cycle. Consequently, any sensitive or offensive content reflects the characteristics of the source material rather than our endorsement of it. Our objective is to investigate and understand political bias in LLMs so that strategies can be developed to prevent such content from disproportionately appearing in real-world deployments. Additionally, we conduct a human review of the news article dataset to remove any identifiable information about article authors.
@@ Appendix A Mathematical Formulation of WMLE
Appendix B Details on Fine-tuning for LLMs
@@
Appendix C Details on Model Training for Political Bias Metric
We experiment with multiple transformer-based models, e.g. BERT (Devlin et al., 2018) and RoBERTa (Liu et al., 2019), to select the best-performing model based on macro F1 score. Each model is fine-tuned using the HuggingFace ‘Trainer‘ class, with a learning rate of $2\times 10^{-5}$, a batch size of 16, and 5 training epochs. We employ a cross-entropy loss function for multi-class classification. Tokenization is performed using each model’s respective tokenizer with a maximum sequence length of 512 tokens. To mitigate overfitting, weight decay of 0.01 is applied during training. The model checkpoints are saved after each epoch, and the best model is selected based on macro F1 score evaluated on the validation set. We use a weighted random sampler during training to ensure balanced class representation. We evaluate the models using the macro F1 score to account for the multi-class nature of the task, ensuring that performance is balanced across all bias categories. The final evaluation is conducted on the held-out test set. Additionally, we report the loss, runtime, and sample processing rates for completeness.
Appendix D Percentage of Center (Left) Biased Articles
Figure 7: This figure shows the percentage of center-leaning articles across multiple generations, comparing the baseline setup (’Synthetic’) with three mitigation strategies.
Figure 8: This figure shows the percentage of left-leaning articles across multiple generations, comparing the baseline setup (’Synthetic’) with three mitigation strategies.
Appendix E Qualitative Bias Analysis Framework and Example of Bias Amplification Across Generations
We employed qualitative methods to confirm our findings in media bias. Specifically, we utilized a media bias identification framework grounded in foundational works such as Entman’s framing theory (Entman, 1993) and other research on media bias detection (Rodrigo-Ginés et al., 2024; Groeling, 2013). This framework provides a robust lens to evaluate political biases in the framing and language use of media texts. Given the nature of our data—text exclusive of visual or contextual cues like formatting—certain types of media bias commonly seen in formatted articles or televised programs (e.g., visual bias or tone) may not apply. Therefore, our focus was on the two key aspects of political bias that are particularly relevant in textual analysis:
@@ Appendix E Qualitative Bias Analysis Framework and Example of Bias Amplification Across Generations
    Analysis of Loaded Language Bias:
@@
Appendix F Distribution of Text Quality Index
Figure 9: This figure presents the distribution of the text quality index across multiple generations for the baseline setup.
Appendix G Average Perplexity Across Generations
Figure 10: This figure shows the average perplexity for synthetic articles , along with 95% confidence intervals, across multiple generations for the baseline setup.
Appendix H Example of Quality Deterioration Across Generations
@@
H.1 Example of Deterministic Setup
@@
  • Generation 4: in Iraq and Syria (ISIS) group Read more The Iraqi army has been fighting the Islamic State since the group seized large swaths of territory in Iraq and Syria (ISIS) group in 2014. The Iraqi army has been fighting the Islamic State the Iraqi army. The move comes as the U.S. the Iraqi the the the the holiest places in the world. Text Quality Index: 2.01; Perplexity: 3.17
@@
Appendix I Mathematical Details for the Statistical Tests
@@
We will explain how the relationship between changes in neuron weight and changes in bias performance (generation quality) can be statistically tested.
@@
Appendix J Literature Review of Model Collapse